OpenStudy (anonymous):

Find the slope of the tangent line @ x=4 (7/x)- sq. root of x

OpenStudy (anonymous):

The derivative of \[\frac{7}{x}-\sqrt{x} \]evaluated at x=4 is\[ -\frac{7}{x^2}-\frac{1}{2 \sqrt{x}}\text{=}-\frac{7}{4^2}-\frac{1}{2 \sqrt{4}}\text{=}-\frac{11}{16}\]Refer to the attached plot.